Fermi National Laboratory


Accelerator Update

Friday August 13
The Day shift began with Tevatron (Tev) store #3725 colliding, AntiProton Source (Pbar) stacking, and with beam to MiniBoone and Switchyard 120. Recycler (RR) was also storing Pbars.

At 10:55 AM the Antiproton Source (Pbar) stopped stacking and went to reverse protons for studies. Switchyard also tested their autotune program at this time.

Pbar studies were completed at 6:03 PM, Pbar then returned to stacking.

At 9:46 PM power had to be dumped to the Kautz Road Substation because of arcing on the feed to the phase B metering transformer. This affects power to the Tev, Main Injector (MI) and the RR. The Tev store had to be terminated. Feeder 46A (Cryo feeder) was back fed to the Master substation, along with the conventional power for the Main Injector. We are unable to run beam to the Main Injector or the Tev until this metering transformer is replaced by ComEd.

Saturday August 14
The Midnight shift began with no beam to MI, Tev, MiniBoone, and Switchyard because of the Kautz Road Substation failure. Pbar is maintaining its stack of 117 e10.

At 1:23 PM Linac and Booster were brought up for Booster studies.

Sunday August 15
The Midnight shift began with no beam to MI, Tev, MiniBoone, and Switchyard because of the Kautz Road Substation failure. Pbar is maintaining its stack of 117 e10.

At 7:41 PM the Kautz Road Substation once again has power, the metering transformer had to be replaced, tested, filled with SF6, and sit for 12 hours before it could be powered. Feeder 46A and conventional power has been switched back to the Kautz Road Substation.

MI and Tev recovery has begun.

Tev cryo was recovered at 9:00 PM.

MI ramp was established at 9:56 PM

Tev ramp was established at 10:42 PM

Monday August 16
The Midnight shift began with recovery of MI, and Tev in progress. MiniBoone, and Switchyard were waiting for beam. Pbar is maintaining its stack of 117 e10.

At 4:31 AM we had beam in the MI.

At 5:38 AM we were stacking Pbars once again.

We have beam in the Tev at 6:36 AM.

Tev is in a wet squeeze at 7:45 AM.

Today's Plans
Complete the recovery and put a store in the Tev.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/16/2004   email Fermilab